How much is the vapor pressure
lowered in an aqueous solution
that has a solute mole fraction of
0.1000? The vapor pressure of
the solvent is 25.756 mmHg at
25oC
1
Expert's answer
2020-04-20T15:06:00-0400

Solution:

χsolvent = 1.0000 - 0.1000 = 0.9000

Use Raoult's Law:

Psolution = (χsolvent) (P°solvent)

x = (0.900) (25.756)

x = 23.18 mmHg (to four sig figs)